### v5

# RefTon: Reference person shot assist virtual Try-on

Abstract
We introduce RefTon, a flux-based person-to-person virtual try-on framework that enhances garment realism through unpaired visual references. Unlike conventional approaches that rely on complex auxiliary inputs such as body parsing and warped mask or require finely designed extract branches to process various input conditions, RefTon streamlines the process by directly generating try-on results from a source image and a target garment, without the need for structural guidance or auxiliary components to handle diverse inputs. Moreover, inspired by human clothing selection behavior, RefTon leverages additional reference images (the target garment worn on different individuals) to provide powerful guidance for refining texture alignment and maintaining the garment details. To enable this capability, we built a dataset containing unpaired reference images for training. Extensive experiments on public benchmarks demonstrate that RefTon achieves competitive or superior performance compared to state-of-the-art methods, while maintaining a simple and efficient person-to-person design.
Figure 1: In-the-wild try-on results produced by our RefTON model. The first row demonstrates our mask-free try-on capability, where the garment is transferred directly to the target person. The second row shows our additional-reference try-on mode, in which extra visual references are incorporated to enhance structural accuracy, texture fidelity, and overall realism.
11footnotetext: Equal contribution.22footnotetext: Corresponding author
1 Introduction
The Virtual Try-On (ViTON) model aims to generate photo-realistic images of a person wearing target clothing, a tool crucial for applications in online retail and personalized fashion systems. ViTON methods are broadly categorized into Generative Adversarial Networks (GANs) [goodfellow2020generative] and Diffusion Models [Ho2020DenoisingDP, Rombach2021HighResolutionIS]. Early ViTON research relied on GANs [choi2021viton, han2018viton, wang2018toward], which typically employed warping modules to deform clothing for alignment with the human body, followed by fusion to achieve visual harmony. However, GAN-based approaches frequently generate unrealistic artifacts, particularly when dealing with complex clothing textures or challenging human poses. Recently, methods based on latent diffusion models (LDMs) [chen2024wear, xu2025ootdiffusion] have gained traction, significantly enhancing clothing warping and addressing structural arrangement and texture preservation during denoising [kim2024stableviton, zhu2023tryondiffusion, xu2024magicanimate, choi2024improving]. Despite these advances, current diffusion-based ViTON technologies still generally rely on extensive auxiliary conditions, such as clothing region masks, garment masks, human poses, key points, or multi-modal inputs like text prompts [xu2025ootdiffusion, kim2024promptdresser, chong2024catvton, chong2025catv2ton, yang2025omnivton].
Despite the remarkable progress of prior virtual try-on approaches, they are still constrained by two critical limitations that hinder the authenticity of the try-on results and broader applicability: First, these approaches rely on multiple external models and internal modules, such as pose estimators [guler2018densepose, toshev2014deeppose, openpose, cao2017realtime, wei2016cpm], human parse models [li2020self, dong2014towards], segmentation models [kirillov2023segment, ravi2024sam2], to process different conditions, which compromises the practicality. To process diverse inputs, additional modules are integrated into the model, which consequently increases the overall framework complexity. Moreover, in practical applications, the quality of conditional inputs—such as the cloth mask—has a substantial impact on the quality of the final try-on results; Second, many aspects of clothing, such as style, texture, and detailed design, cannot be fully perceived from the garment image alone; instead, it is more important to consider the overall appearance when a model wears the garment. Therefore, in real-world try-on scenarios, such as online shopping, users are typically more interested in model images rather than the garment itself. They tend to see how the target garment looks when worn on a real person, rather than relying solely on the isolated garment image as a reference. For example, as illustrated in Fig. 2, at the garment in the first row, it is difficult to tell whether it is a green translucent fabric or a light green opaque one, whereas the reference image clearly reveals its green translucent material. We cannot accurately identify transparent materials or intricate designs, such as lace collars, solely from cloth images. In contrast, the reference images of human models wearing the garments reveal such details. However, existing virtual try-on methods do not support such references due to the lack of corresponding reference data in public datasets [morelli2022dress, choi2021viton, fang2024vivid, zheng2019virtually, shen2025imagdressing, he2024wildvidfit, meng2025hf].
Figure 2: Effect of reference images in virtual try-on. From left to right: inputs (agnostic, cloth, reference), results without reference images, results with reference images, and ground truth. Using reference images in both training and inference improves visual fidelity and preserves fine details (e.g., transparent materials and lace collars). Zoom in for a better comparison.
Based on the above observation, we propose RefTon, a flux-based person-to-person virtual try-on framework that achieves strong performance without relying on any external models or auxiliary components, while being further enhanced by additional reference images that offer more accurate and context-aware guidance for the try-on model. First, to ensure the best performance, we adopt the powerful image editing model Flux-kontext as our base and apply adaptation on its position index for RoPE [su2024roformer] to make it suitable for multi-condition/resolution virtual try-on inputs. Similar to [chong2024catvton], RefTon eliminates the need for auxiliary inputs such as segmentation masks using a two-stage training strategy, allowing for simple inference with only a source image and the target garment as inputs. Second, we introduce the use of images of a different person clothed in the target garment as the visual references, like the model images in online shopping, which better reflect users’ real-world behavior when choosing clothes and enable the preservation of fine garment details that existing methods cannot achieve. To achieve these objectives, we propose a reference data generation pipeline, by which we construct a dataset with supplementary reference images and use unpaired person-cloth samples to train our own model to utilize reference images as additional visual guidance. These improvements empower RefTon to achieve both a simplified model structure and a streamlined inference process, while simultaneously delivering superior generation results.
In summary, the main contributions are as follows:
  • We propose incorporating additional reference images into the virtual try-on pipeline. This significantly enhances the authenticity and visual quality of the try-on results, achieving State-of-the-Art performance in preserving fine garment design details.
  • We designed a reference data generation framework to create the necessary reference images for both the clothing and target ground truth samples. Based on this pipeline, we built the VFR dataset upon existing benchmarks (e.g., VITON-HD, DressCode, ViViD), providing a robust new resource to improve the practicality and evaluation of virtual try-on models.
  • We present an adaptation of the Flux-Kontext I2I model with a modified Rescaled Position Indexing mechanism to support flexible multi-conditional and multi-resolution inputs, along with a two-stage training strategy for virtual try-on. Our framework enables integration of varying numbers and types of reference images and effectively supports mask-based and person-to-person try-on within a single model. It achieves state-of-the-art performance and demonstrates strong generalization to in-the-wild person–clothing scenarios.
2 Related Works
2.1 Generative Model via Flow Matching
Generative modeling has rapidly progressed with diffusion models (DMs) [sohl2015deep], score-based generative models (SGMs) [song2020score], and flow-based methods. Recent work has explored controllable generation, layout-conditioned synthesis, unified planning-generation frameworks [cheng2024hico, he2025plangen], as well as diffusion transformer variants for multilingual generation, efficient sampling, and fine-grained control [Liu_2025, ma2025nami, liu2025nanocontrollightweightframeworkprecise, gong2025cta]. Flow-based methods [dinh2016density, dinh2014nice, kingma2018glow] have been further advanced to address the inefficiency of continuous normalizing flows (CNFs), which require costly backpropagation through ODE solvers during training [Chen2018NeuralOD]. Flow Matching (FM) [Lipman2022FlowMF] mitigates this limitation by learning a time-dependent vector field that deterministically transports a simple before the data distribution, using a simulation-free objective that avoids numerical integration during training. By directly parameterizing the probability flow, FM achieves competitive or superior sample quality compared to diffusion models with significantly fewer sampling steps. Our method builds upon the Flux-Kontext architecture, where input images are encoded into latent representations, flattened into sequences, and concatenated with Gaussian noise $\bm{\epsilon}$. It is also closely related to recent FLUX-based appearance editing and transfer approaches [zhu2025flux].
2.2 Diffusion-based Virtual Try-on
Diffusion models [Ho2020DenoisingDP, Rombach2021HighResolutionIS] have enabled significant progress in garment or makeup transfer [kim2024stableviton, zhu2023tryondiffusion, choi2021viton, zhu2025flux]. Leveraging the flexibility of Stable Diffusion [morelli2023ladi, kim2024stableviton], prior works exploit text guidance and inpainting for garment synthesis. Extensions such as DiffusionCLIP [kim2022diffusionclip] introduce semantic control via CLIP, while methods like DCI-VTON [gou2023taming] and IDM-VTON [choi2021viton] adopt two-stage pipelines to align and fuse garments, improving structural consistency.
Recent approaches, including CatVTON [chong2024catvton], OmniTry [feng2025omnitry], Any2AnyTryon [guo2025any2anytryon], and OmniVTON [yang2025omnivton], explore person-to-person try-on without explicit masks. However, these methods typically rely on additional conditions (e.g., pose) or fail to unify mask-based and mask-free settings. Moreover, P2P pipelines such as TryOffDiff [velioglu2025enhancing] and ViTON-GUN [ViTON-GUN] adopt a “try-off–then–try-on” strategy, which introduces error accumulation and loss of garment details. In contrast, our method directly leverages the reference image, avoiding the try-off stage and better preserving garment structure and material fidelity. In summary, existing diffusion-based virtual try-on methods either rely on heavy auxiliary annotations or lack support for clothed reference images. To address these limitations, we propose RefTON, which adapts the virtual try-on task to the Flux-Kontext framework, enabling end-to-end, reference-guided generation while supporting both mask-based inpainting and mask-free editing.
Figure 3: Overview of our two-stage training strategy. (a) Stage 1 follows mask-based try-on paradigms to generate person images wearing random garments from masked inputs, providing training data for Stage 2. (b) Stage 2 uses synthesized person images, along with the target garment and optional reference images, to train a person-to-person virtual try-on model that directly fits the garment onto the person.
3 Method
3.1 Preliminary
RefTon is built upon DiT [peebles2023scalable], a scalable Transformer architecture for diffusion-based generation. Images are encoded into a latent space via an autoencoder [kingma2013auto] and then patched into tokens [dosovitskiy2020image]. The diffusion process [Ho2020DenoisingDP] operates on these tokens, with the Transformer consuming noisy tokens and predicting their denoised results.
We consider the problem of generating images under the condition $\bm{y}$, which may represent garment images, semantic maps, human pose, or other modality-specific control signals. Let $\bm{x}$ denote the latent image representation obtained from a VAE encoder. The goal of Flux.1 [flux2024, labs2025flux1kontextflowmatching] is to approximate the conditional distribution $p(\bm{x}\mid\bm{y})$ by learning a time-dependent velocity field $\bm{v}(\bm{x},\bm{y},t)$ that transports a sample from a simple prior $\mathcal{N}(\bm{x};\bm{0},\bm{I})$ at $t=0$ to the data distribution $p_{\text{data}}(\bm{x}|\bm{y})$ at $t=1$. The dynamics of the conditional probability density $p(\bm{x}|\bm{y},t)$ over time $t$ are governed by the continuity equation:
$\frac{\partial}{\partial t}p(\bm{x}|\bm{y},t)$ $=-\nabla_{\bm{x}}\cdot\big(\bm{v}(\bm{x},\bm{y},t)\cdot p(\bm{x}|\bm{y},t)\big),\quad$ (1)
$\bm{x}_{t=0}\sim\mathcal{N}(\bm{x};\bm{0},\bm{I}),\ \bm{x}_{t=1}\sim p_{\text{data}}.$
To estimate $\bm{v}(\bm{x},\bm{y},t)$, we train a diffusion-transformer backbone to approximate the neural velocity field $\bm{v}_{\bm{\theta}}$ using the conditional flow matching objective [Lipman2022FlowMF, liu2022flow]:
$\mathcal{L}_{\bm{\theta}}$ $=\mathbb{E}_{t,\bm{x}_{i},\bm{\epsilon},\bm{y}_{i}}\Big[\big\|\bm{v}_{\bm{\theta}}(\bm{x},\bm{y}_{i},t)-(\bm{x}_{i}-\bm{\epsilon})\big\|_{2}^{2}\Big],$ (2)
$\bm{x}=(1-t)\,\bm{x}_{i}+t\,\bm{\epsilon},$
where $t\sim\mathcal{U}(0,1)$, $\bm{x}_{i}\sim\mathcal{X}_{\mathrm{train}}$, and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. This training objective encourages the model to learn a velocity field $\bm{v}_{\bm{\theta}}(\bm{x},\bm{y},t)$ that consistently guides the noisy samples toward the data distribution conditioned on $\bm{y}$, following the probability flow ODE starting from the Gaussian prior:
$d\bm{x}=\bm{v}(\bm{x},\bm{y},t)dt,$ (3)
enabling controllable image synthesis at inference time.
In the virtual try-on setting, let $\bm{x}_{i}$ denote the image of a person wearing the target cloth, and let $\bm{y}_{i}$ represent a collection of conditional inputs, including the cloth-agnostic image $\bm{a}_{i}$, the target cloth $\bm{c}_{i}$, the visual references $\bm{r}_{i}$, and others. Formally, we write $\bm{y}_{i}=[\bm{a}_{i},\bm{c}_{i},\dots].$ The objective is to progressively transform a Gaussian noise sample $\bm{\epsilon}$ into the target image $\bm{x}_{i}$ guided by conditions $\bm{y}_{i}$.
3.2 Person To Person Virtual Try-on Model with Two Stage Training
Our goal is to dress the person directly with the target garment, without relying on auxiliary conditions such as DensePose [guler2018densepose] or segmentation masks. To achieve this, we train a diffusion model on clothing-person pairs $(\bm{c}_{i},\bm{\bar{p}}_{i,\bm{c}_{j}})$ to generate the target image $\bm{p}_{i,\bm{c}_{i}}$, where the person wears the target garment, as shown in Fig. 3(b). This setup requires unpaired triplets $\bm{\bar{p}}_{i,\bm{c}_{j}},\bm{c}_{i},\bm{p}_{i,\bm{c}_{i}}]$, where $\bm{c}_{j}\neq\bm{c}_{i}$. However, existing open-source benchmarks typically provide only paired data $\bm{c}_{i},\bm{p}_{i,\bm{c}_{i}}]$. Following the two-stage training strategy of CATVTON [chong2024catvton], we adopt a similar pipeline for our try-on model and further exploit richer conditions, including agnostic masks and DensePose, to improve unpaired image generation. Specifically, in the first stage of RefTon training, we synthesize unpaired person images $\bm{\bar{p}}_{i,\bm{c}_{j}}$ using a mask-based try-on model.
As illustrated in Fig. 3(a), we train a virtual try-on model using agnostic person images $\bm{a}_{i}$, clothing images $\bm{c}_{i}$, densepose maps $\bm{d}_{i}$, warp masks $\bm{m}_{i}$ as inputs. During inference, a random unpaired garment $\bm{c}_{j}$ is selected to generate the corresponding synthesized person image $\bm{\bar{p}}_{i,c_{j}}$. To ensure the quality of the synthesized person image $\bm{\bar{p}}_{i,c_{j}}$, the agnostic–cloth pairs $\bm{a}_{i},\bm{c}_{i}]$ used for training must belong to the same garment category (e.g., if $\bm{a}_{i}$ is from the “dresses” subset, the selected garment $\bm{c}_{j}$ should also come from “dresses” rather than “upper body” or “lower body”). Otherwise, the generated person image may appear unrealistic due to mismatches between the clothing mask region and the target garment (e.g., fitting a skirt onto the upper body or a shirt onto the lower body).
After obtaining the unpaired person images, we train the person-to-person model with either the agnostic image $\bm{a}_{i}$ or the generated person image $\bm{\bar{p}}_{i,c_{j}}$ from the first stage, sampled with equal $50\%$ probability. Instead of training a try-on model from scratch, we freeze the encoder and decoder of Flux-kontext [labs2025flux1kontextflowmatching] and fine-tune only the transformer blocks using Low-Rank Adaptation [hu2022lora]. The model is optimized with the flow-matching objective in equation. 2. In addition, to help the model capture garment appearance from another person, we provide an extra reference image $\bm{r}_{i}$ with a probability of $25\%$, as illustrated in Fig. 3(b).
Figure 4: Rescaled three-channel position index. (a) The first channel encodes conditional inputs, while the second and third capture spatial positions for varying resolutions. (b) The concatenated indices are used to generate RoPERoPE, which is applied to the latent sequence in the attention mechanism.
3.3 Multi-input Training and Adaptation
The latent embeddings are concatenated into the sequence $\bm{\epsilon}_{i},\bm{a}_{i},\bm{c}_{i}]$ after image encoding in vanilla Flux-kontext, and a index generator produces a position index of shape $[L,3]$, which is further converted into Rotary Position Embeddings (RoPE) for self-attention. Its first channel is a binary mask that distinguishes Gaussian noise $\bm{\epsilon}_{i}$ from conditional image/text inputs, while the second and third channels encode horizontal and vertical coordinates, respectively (Fig. 4). However, this binary design is insufficient for our setting, where the model must handle multiple heterogeneous image conditions. Therefore, we extend the first channel from a binary flag to discrete condition labels, allowing the transformer to distinguish different input types such as person, garment, and reference images.
Specifically, our model follows the Flux-kontext to encode each image condition independently into latent patches, and concatenates to form a sequence such as $\bm{a}_{i},\bm{c}_{i},\bm{r}_{i},\ldots]$, which is then concatenated with the noise latent $\bm{\epsilon}_{i}$ along the sequence dimension. For each condition, we generate an individual three-channel position index: the first channel indicates the condition identity, and the second and third channels store integer spatial coordinates rescaled by the resolution ratio between the target image and the corresponding condition image, which preserves spatial alignment across different resolutions. This Rescaled Position Index design enables flexible integration of multi-type and multi-resolution image conditions within a unified DiT framework. Similar indexing strategies have been explored in prior work [guo2025any2anytryon], where positional indices are constructed on a concatenated image canvas, placing different conditions in a shared spatial layout. In contrast, our method assigns positional indices to each condition independently rather than generating them from a pixel-space concatenated canvas, resulting in a more flexible formulation for multi-resolution inputs. The modified DiT position indexing scheme is illustrated in Fig. 4.
Figure 5: Reference image generation pipeline. Given a target image of a European woman wearing the target upper-body garment, Qwen2.5-VL [bai2025qwen2] produces an appearance description (e.g., ”a European woman with brown straight hair”) and an opposite description (e.g., ”an African woman with curly blonde hair”). In (a), the opposite description is combined with the action and non-target clothing attributes to form the positive prompt, while the appearance description is used as the negative prompt. In (b), the image and prompts are fed into Flux-kontext to generate reference images of different individuals wearing the target garment.
3.4 Virtual-Tryon Generation with Extra Visual Reference
For both humans and generative virtual try-on models, garment images and conditions extracted by external models alone are insufficient to capture the realistic visual effect of wearing a target garment, as illustrated in Fig. 2. A garment’s style, texture, and fine design details are more faithfully presented when worn by another person than when shown in isolation. Therefore, we introduce reference person images $\bm{r}_{i}$, in which another person wears the target garment, to provide more intuitive visual guidance during virtual try-on generation. To obtain such references, we construct pairs of the form $\bm{c}_{i},\bm{r}_{i}]$, representing ”different persons wearing the target garment.” Since these reference images $\bm{r}_{i}$ are unavailable in existing open-source virtual try-on datasets, we develop a reference data generation pipeline to synthesize them. This pipeline augments current datasets with supplementary references and enables model training with additional visual guidance.
Existing open-source datasets like VITON-HD [choi2021viton], DressCode [morelli2022dress], and IGPairs [shen2025imagdressing], lack unpaired reference images $\bm{r}_{i}$ showing different persons wearing the target garment. To overcome this limitation, we employ editing models to synthesize such reference images. To serve as accurate and informative visual guidance, the generated reference data $\bm{r}_{i}$ should satisfy the following requirements:
  • Preserve the target garment faithfully. The target garment’s color, texture, and design must remain unchanged to ensure an accurate reference.
  • Introduce diversity in the person’s appearance. The person wearing the target garment in the reference image should be different from the target person wearing the same garment. Otherwise, the model may learn shortcuts and overfit to the target image. This diversity can be achieved by altering hairstyle, hair color, skin tone, body pose, or facial expression.
  • Vary the non-target garments to provide outfit diversity. While the target garment remains unchanged, other garments should be modified. For example, if the target garment is an upper-body item, the reference image should retain the same upper-body garment while altering the lower-body clothing, shoes, or accessories.
Since the reference image is edited from the target image $\bm{p}_{i,\bm{c}_{i}}$ in existing datasets, the first requirement ensures faithful garment preservation, while the second and third promote diversity in non-target regions to prevent overfitting and prevent the model from taking the shortcut of directly copying the target image in the generated try-on results.
As illustrated in the pipeline, we use Flux-kontext [labs2025flux1kontextflowmatching] to synthesize reference-person images. Given a target image and carefully designed text prompts, Flux-kontext generates images that faithfully preserve the target garment’s color, texture, and design while varying the wearer’s appearance and non-target clothing. Specifically, we first feed the target image and a textual instruction into Qwen2.5-VL [bai2025qwen2] to obtain a detailed description of the person’s appearance, together with an opposite description. We then collect a set of non-target garments and action descriptions to introduce diversity in clothing and pose; examples of these garments and action items are provided in Appendix A. Finally, we concatenate the opposite appearance description with the action and non-target garment descriptions to form the positive prompt, while the original appearance description is used as the negative prompt. These prompts are then fed into Flux-kontext to synthesize the reference images.
We supplement existing virtual try-on benchmarks including VITON-HD [han2018viton], DressCode [morelli2022dress], ViViD [fang2024vivid], FashionTryOn [zheng2019virtually] and IGPairs [shen2025imagdressing] by generating corresponding reference pairs for each target garment image $\bm{c}_{i}$, forming data pairs $\bm{c}_{i},\bm{r}_{i}]$ that are used for both model training and evaluation.Some open-source datasets, such as FashionTryOn [zheng2019virtually] and IGPairs [shen2025imagdressing], contain numerous duplicated or low-quality samples. We compare the CLIP features of images to filter out redundant samples and employ Qwen2.5-VL to identify distorted or unclear images, as well as images where the person faces away from the camera, ensuring overall data quality before generating the final reference set via our data generation pipeline. Finally, we enrich open-source virtual try-on datasets with additional visual references $\bm{r}_{i}$ and person images $\bm{\bar{p}}_{i,c_{j}}$, and combine them to form our own dataset, named Virtual Fitting with Reference (VFR), for training our RefTon model. The detailed data collection, filtering procedures, and visualization of the samples are provided in Appendix A.
Figure 6: Qualitative comparison on the VITON-HD dataset. “reference” denotes using additional reference $\bm{r}_{i}$ for the inference.
Figure 7: Qualitative comparison on the DressCode dataset., and the model is trained following the pipeline in Fig. 3 (b). “reference” denotes the reference $\bm{r}_{i}$ image is used during the inference.
4 Experiments
We mainly evaluate our method on two public benchmarks, DressCode [morelli2022dress] and VITON-HD [choi2021viton], both containing images with a resolution of $1024\times 768$. The VITON-HD comprises 13,670 upper-body image pairs of women, split into 11,647 pairs for training and 2,032 pairs for testing. The DressCode includes three subsets—upper body, lower body, and dresses—with 48,392 training and 5,400 testing pairs. Since DressCode does not provide wrapped cloth masks or agnostic images $\bm{a}_{i}$, we generate them using the mask generation tool from CatVTON [chong2024catvton]. We further conducted experiments on the StreetTryOn [cui2023street-tryon] and included comparisons with the methods mentioned to demonstrate the model’s P2P and in-the-wild capabilities. In the P2P evaluation, we did not use the cloth-only image as input and only used the person image as a reference.
During training, we fine-tune our model on the flux-kontext backbone using the Low-Rank Adaptation (LoRA) technique, with a rank of $64$ and $\alpha=128$, optimized by AdamW. For quantitative evaluation, all generated images are resized to $512\times 384$ to ensure a fair comparison with previous methods. In a single-dataset experiment, the model is trained independently for 20,000 steps on VITON-HD and 48,000 on DressCode with a batch size of 128, using 8 NVIDIA H100 GPUs. To further enhance generalization and robustness to in-the-wild inputs, we train an additional model on the mixed VFR dataset introduced in Sec. 3.4. We report cross-dataset evaluation results of our RefTon model at a resolution of $1024\times 768$ on both VITON-HD and DressCode, and demonstrate its in-the-wild performance using images collected by ourselves.
4.1 Quantative result
We provide the numerical results of our model on the VITON-HD and DressCode datasets, distinguishing between paired and unpaired try-on settings. For paired try-on settings with ground truth in test datasets, we utilize four metrics to assess the similarity between the synthesized images and their corresponding authentic images: the Structural Similarity Index (SSIM) [nilsson2020understanding], Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], Fréchet Inception Distance (FID)[Seitzer2020FID], and Kernel Inception Distance (KID)[binkowski2018demystifying]. For unpaired settings, where we measure the distributional similarity between the synthesized and real samples, we specifically rely on the Fréchet Inception Distance (FID) and Kernel Inception Distance (KID). As shown in Table 1.
Table 1: Quantitative comparison on VITON-HD [choi2021viton] and DressCode [morelli2022dress]. Best and second-best results are shown in bold and underline, respectively. “+R” denotes the use of reference images, “MF” indicates mask-free inputs, and “–” denotes missing values. Subscripts $p$ and $u$ denote the paired and unpaired settings, respectively.
Method Input VITON-HD DressCode
[…]
CAT-DM [zeng2024cat] 0.080 0.877 5.60 0.83 8.93 1.37
IDM-VTON [choi2024improving] 0.102 0.870 6.29 0.062 0.920 8.64 0.90
OOTDiffusion [xu2025ootdiffusion] 0.071 0.878 8.81 0.82 0.045 0.927 4.20 0.37
CatVTON [chong2024catvton] 0.057 0.870 5.43 0.41 9.02 1.09 0.046 0.892 3.99 0.82 6.14 1.40
CatVT2ON [chong2025catv2ton] 0.057 0.890 8.10 2.25 11.22 2.99 0.037 0.922 5.72 2.34 8.63 3.84
OmniVTON [yang2025omnivton] 0.145 0.832 7.76 9.62 0.119 0.865 5.34 6.45
[…]
PromptDresser [kim2024promptdresser] 0.112 0.869 8.54 0.67
RefTON (Ours) 0.057 0.873 5.45 0.82 8.58 1.06 0.037 0.912 3.48 1.20 5.31 1.36
RefTON+R (Ours) 0.049 0.879 4.69 0.68 8.43 0.91 0.031 0.918 2.94 0.95 5.07 1.15
Mask-Free setting
CatVTON(Mask-Free) [chong2024catvton] 0.061 0.870 5.89 0.51 9.29 1.17 0.045 0.902 4.78 1.30 7.40 2.62
Any2AnyTryon [guo2025any2anytryon] 0.088 0.839 6.93 0.74 8.97 0.98
TryOffDiff [velioglu2025enhancing] 11.9 2.60 7.90 2.70
RefTON/MF (Ours) 0.061 0.866 5.98 1.04 8.40 0.81 0.041 0.901 3.84 1.33 5.00 1.17
RefTON+R/MF(Ours) 0.053 0.872 5.11 0.82 8.32 0.78 0.035 0.906 3.34 1.15 5.02 1.28
Our method (RefTon) consistently performs better than prior baselines, demonstrating higher try-on fidelity and strong alignment with the target person’s pose. With the addition of reference images (“+R”), the quality and detail consistency of the try-on results are further improved compared with the results without reference images, establishing new state-of-the-art results across multiple metrics. Notably, even in the mask-free setting—without agnostic masks or auxiliary inputs—our method maintains garment style correctness and pose consistency, while reaching accuracy on par with or superior to baseline methods, highlighting its robustness and practicality.
Table 1 summarizes the quantitative evaluation on the DressCode dataset. Our method (RefTon) outperforms all baselines, delivering higher try-on quality and achieving strong consistency with the target person’s pose and body structure. Integrating reference images (“+R”) further enhances the results, establishing a new state-of-the-art. Importantly, even in the mask-free setting—without agnostic masks or additional inputs—our method correctly preserves garment styles (e.g., clothing length and design) and maintains high pose alignment, while achieving accuracy comparable to or surpassing prior baselines, demonstrating robustness and practicality. Additional quantitative results on the DressCode subset are provided in Appendix B in detail.
Results in Table 2 show that our method achieves state-of-the-art performance on the StreetTryOn benchmark. Notably, our model is neither trained on the StreetTryOn domain nor designed to handle the absence of garment images during training. These results demonstrate its strong generalization ability and flexibility, and further confirm that it can directly extract garment information from the reference image for virtual try-on.
Table 2: Quantitative comparisons on the StreetTryOn. Sh, St, P, denotes the shop, street, and person(model) image, respectively.
Method Required Input Sh-to-St P-to-P P-to-St St-to-St
Mask Pose Text $\downarrow FID\downarrow$ $\downarrow FID\downarrow$
StreetTryOn 34.054 12.185 34.191 33.039
OmniVTON 33.919 8.983 33.450 23.470
RefTON 28.991 8.870 25.429 16.452
4.2 Qualitative comparison
Previous methods can fit garments onto a person, but often fail to preserve wearing-critical attributes such as cut, style, and fine details. By contrast, our method achieves improved visual fidelity among all baselines. As shown in Fig. 6, on VITON, RefTon realistically renders challenging materials such as hollow and semi-transparent fabrics even without reference images. For example, it faithfully preserves the perforated structures and transparency of lace garments, while baselines often produce solid textures or spurious dotted artifacts. Our method also better maintains garment patterns, keeping printed letters and logos clear and consistent with the input. Adding reference images further improves generation quality, validating the effectiveness of our reference assist framework. Even without agnostic masks, directly transferring clothing $\bm{c}_{i}$ onto person images $\bm{\bar{p}}_{i,\bm{c}_{j}}$ still allows our method to outperform most baselines. More results on VITON-HD, DressCode, and in-the-wild examples are provided in Appendix C.
Table 3: Cross-dataset comparison with OOTDiffusion [xu2025ootdiffusion] on VITON-HD and DressCode under both paired and unpaired settings. For the mask-free setting, we report only the unpaired results, as the paired setting is not meaningful when no garment–person alignment is enforced.
Methods VITON-HD (Paired) DressCode (Paired)
[…]
OOTDiffusion [xu2025ootdiffusion] 0.839 11.22 2.72 0.123 0.915 11.96 1.21 0.061
RefTON (Ours) 0.851 6.23 0.80 0.072 0.896 3.70 1.13 0.045
RefTON+R (Ours) 0.859 5.13 0.62 0.060 0.903 3.14 0.97 0.038
Methods VITON-HD (Unpaired) DressCode (Unpaired)
$downarrow FIDu\downarrow KIDu\downarrow$
RefTON (Ours) 9.11 1.08 5.22 1.20
RefTON+R (Ours) 8.59 0.87 5.03 1.11
RefTON/MF (Ours) 8.88 0.82 5.03 1.23
RefTON+R/MF (Ours) 8.39 0.65 4.87 1.10
Our model also performs well on a benchmark with a wider variety of clothing types. As shown in Fig. 7, we evaluate our method on the DressCode dataset, where garments are categorized into upper body, lower body, and dresses. Our approach produces more faithful and natural try-on results compared to previous methods. In particular, it renders reflective materials, such as leather and metallic fabrics, with superior realism, avoiding the over-smoothing or distortion artifacts commonly observed in other methods. Moreover, even without agnostic masks, our model can still perform consistent try-on guided by garment style, accurately preserving the length, structure, and overall design without introducing mismatched or inconsistent shapes.
4.3 Training on VRF Dataset and Evaluation
To evaluate the effectiveness of our person-to-person virtual try-on framework and data construction pipeline, we build a mixed dataset, Mixed-Virtual-Ref, by combining samples from DressCode, VITON-HD, FashionTryOn [xiao2025omnigen], ViViD [fang2024vivid], and IGPairs [choi2024improving]. We use Qwen2.5-VL to filter low-quality images, generate agnostic masks via [chong2024catvton], and obtain reference and unpaired person images through our data generation process. In total, 103,936 image pairs are collected for training, using the same hyperparameters as in the DressCode and VITON-HD experiments.
We evaluate the model on the DressCode and VITON-HD test sets, with quantitative results shown in Table 3. Our method (RefTon) achieves consistently superior or comparable performance to OOTDiffusion [xu2025ootdiffusion] across both benchmarks and both paired and unpaired settings. Notably, despite not being trained on DressCode or VITON-HD individually, the mixed-dataset model surpasses dataset-specific baselines on most metrics (e.g., FID and LPIPS), demonstrating strong cross-dataset generalization and the robustness of our data construction strategy.
4.4 Ablation Study
Figure 8: Qualitative results of the ablation study across different settings. “Ref.” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.
We conduct an ablation study to examine our model under four settings(w/&w/o mask,w/&w/o Ref.). As shown in Table 1, our model maintains consistently strong performance across all settings. Introducing a reference image yields clear improvements in both mask-based and mask-free modes, while moving from mask-based to mask-free inputs causes only mild metric fluctuations, confirming the model’s stable robustness without masks.
Fig. 8 further provides qualitative comparisons. The first two rows show that reference images help the model correctly infer garment structures and materials that are ambiguous in the flat images (e.g., hollow textures, semi-transparency). The last two rows illustrate that mask quality heavily affects mask-dependent models: overly aggressive masks remove important items (e.g., handbags), while conservative masks retain unwanted regions (e.g., legs), leading to incorrect garment geometry. In contrast, our mask-free model consistently produces correct outputs regardless of mask or reference conditions, demonstrating that mask-free capability reduces reliance on mask quality and enables more flexible and stable try-on performance.
To validate the rescaled position index, we experiment on VITON-HD with cloth image $\bm{c}_{i}$, dense pose $\bm{d}_{i}$, warp mask $\bm{m}_{i}$, and reference image $\bm{r}_{i}$ as conditions. In the masked and mask-free settings, the target cloth is transferred to the person image and the agnostic image, respectively. We evaluate two condition scales, $0.25\times$ and $0.5\times$, where the person or agnostic image is resized to the target resolution and the conditional inputs are rescaled accordingly. As shown in Table 4, the rescaled position index outperforms the original Flux-Kontext position index on FID and KID under both settings, showing its effectiveness for multi-condition inputs with varying resolutions.
Table 4: Rescaled position index (PI) ablation.. We compare the original Flux-Kontext position index with our rescaled PI at different condition scales; MFMF denotes mask-free inputs. Conditional inputs are resized to $0.5\times$ and $0.25\times$ of the target resolution.
Method 0.5 Scale 0.25 Scale
[…]
w/ o Rescaled PI 5.29 0.84 4.75 0.72 6.08 0.93 5.35 0.76
w/ Rescaled PI 5.09 0.79 4.71 0.69 6.01 0.77 5.37 0.71
5 Conclusion
This paper introduces RefTon, a virtual try-on framework that supports both mask-based and mask-free inference and leverages reference images to guide the try-on process. We extend Flux-Kontext to handle multi-condition inputs of varying resolutions via a rescaled position index. To train RefTon, we propose a reference data generation pipeline integrating Qwen2.5-VL and Flux-Kontext. This design allows RefTon to faithfully preserve translucent fabrics, intricate designs, and fine details, consistently outperforming existing methods both quantitatively and qualitatively, achieving state-of-the-art performance in all settings.
6 Acknowledgments
This research was supported by 360 AI Research. We sincerely thank 360 AI Research for its support of this work.
Supplementary Material
Appendix A Generation of Reference Image
This section provides a detailed description of the process for generating the reference image. Many virtual try-on datasets offer the garment image and the image of the target person wearing the target garment, but they do not include the reference image, which shows the visual effect of the target garment $\bm{c}_{i}$ being worn by another person $\bm{p}_{\bm{c}_{j}}$. The generation of the reference image can be viewed as an image editing task on $\bm{p}_{\bm{c}_{j}}$. As discussed in Section 3.3, the reference image $\bm{r}_{i}$ must satisfy three key requirements.
Firstly, the reference image should faithfully preserve the details of the target garment $\bm{c}_{i}$. This requires the editing model to have strong consistency abilities. The Flux-Kontext model has a robust ability to edit the target region corresponding to the text prompt while keeping unrelated regions—such as the area not related to the garment—unchanged. Moreover, the Flux-Kontext model can perform precise local editing according to the text prompt, in this case, focusing on the person under the garment. Therefore, we choose the popular Flux-Kontext model [36] to edit the input image conditioned on the text prompt. Specifically, we add a sentence such as “keep the {target cloth} cloth unchanged” in the text prompt.
Secondly, the person in the reference image should look significantly different from the person in the target image. During training, we observed that if the reference image is too similar to the target image, the model tends to rely on a ”shortcut”—directly copying from the reference image and ignoring the agnostic/person image $\bm{a}_{i}$$\bm{p}_{i}$ and the cloth $\bm{c}_{i}$. To avoid this, we ensure that the person in the reference image differs from the target image to better showcase the visual effect of the clothing when worn, rather than focusing on the appearance of the person themselves. To achieve this, we utilize the Text-to-Image (T2I) capabilities of the Flux-Kontext model. We extract an accurate description of the person’s appearance in the target image (e.g., ”The model has an East Asian appearance, with light skin, long black hair, and a neutral expression…”) and pass it as the negative prompt to the T2I model. In contrast, we provide an opposite description (e.g., ”The model has an African appearance, with dark skin, short yellow hair, and a cheerful expression…”) as the positive prompt to guide the editing of the target image, as shown in Figure 4(b).
However, extracting descriptions for each image manually is labor-intensive. To address this, we use a vision-language model (VLM) such as Qwen2.5-VL to automatically generate the description and its opposite. Specifically, we pass the target image $\bm{p}_{i,\bm{c}_{i}}$ to the VLM and provide a prompt like “Start with Positive: describe only the model’s race, skin, hair, eyes, and expression, then give the opposite in one sentence with Negative: changing those traits without ’not’ or clothing.” The generated description and its opposite are then fed into the negative and positive prompt encoders of the T2I model to edit the target image, as shown in Figure 4(a). In this way, we automate the generation of the reference image $\bm{r}_{i}$ by editing the target image.
Thirdly, after extracting the description and opposite description of the person’s appearance, we introduce more diversity into the reference image by varying the non-target garments and actions of the human in the target image. This can be achieved through the image editing model by adding descriptions related to non-target garments and actions. We provide a description bank containing candidate descriptions for outfits and actions across three scenarios: the person in the image is wearing a dress, an upper-body garment, or a lower-body garment. This ensures that the description of the editable garment differs from the target garment $\bm{c}_{i}$. Furthermore, the outfit descriptions also include accessories such as glasses, wristwatches, and bracelets to increase diversity. These descriptions, along with the actions and outfit details, are concatenated into positive prompts and passed to the T5 text encoder, as shown in Figure 4.
Fig S9 illustrates selected text prompts from the prompt description bank used for reference image generation. Furthermore, Fig S10 exhibits a selection of the resulting reference data samples.
Figure S9: Sample text prompts from the Outfit and Action Description Bank. To ensure the model edits only the person while preserving the target clothing, we assign different outfits and action description categories to different clothing inputs.
Figure S10: Sample reference images generated by our reference data generation pipeline. The editing model takes the target person’s image as input and synthesizes corresponding reference images, while preserving the garment’s appearance to match the cloth.
Appendix B Further Quantitative Evaluation on High Resolution
Here we provide more results of the generated image. Table S5 summarizes the detailed quantitative evaluation on the DressCode dataset. Our method (RefTON) outperforms all baselines, delivering higher try-on quality and achieving strong consistency with the target person’s pose and body structure. Integrating reference images (“+R”) further enhances the results, establishing a new state-of-the-art. Importantly, even in the mask-free setting—without agnostic masks or additional inputs—our method correctly preserves garment styles (e.g., clothing length and design) and maintains high pose alignment, while achieving accuracy comparable to or surpassing prior baselines, demonstrating robustness and practicality.
Table S5: Quantitative results on three subsets of the DressCode dataset [morelli2022dress]: upper body, lower body, and dresses. The best and second-best results are highlighted in bold and underline, respectively. The symbol “*” denotes results reported in prior work, while “+R” indicates results with reference image input, and “MF” refers to mask-free input images. The subscripts $p$ and $s$ denote specific evaluation metrics for precision and recall, respectively.
Method Upper-body Lower-body Dresses
[…]
CAT-DM [zeng2024cat] 9.85 2.38 12.62 1.89 10.25 1.81 14.83 2.82 10.71 2.02 14.30 3.36
OOTDiffusion  [xu2025ootdiffusion] 11.03 0.29 9.72 0.64 10.65 0.54
PromptDresser [kim2024promptdresser] 11.00 0.74 12.55 1.46 11.09 1.10
RefTON(Ours) 7.62 1.10 11.13 0.98 7.60 1.38 13.07 2.11 7.32 1.30 11.56 1.98
RefTON+R(Ours) 6.39 0.85 11.08 0.87 6.61 1.05 12.56 1.67 6.09 1.16 11.16 1.72
RefTON/MF(Ours) 8.37 1.43 11.20 1.11 8.79 1.51 12.50 1.83 7.36 1.33 10.73 1.41
RefTON+R/MF(Ours) 7.20 1.18 11.53 1.12 7.85 1.21 12.74 2.05 6.24 1.20 10.05 1.30
Table S6: Quantitative comparison across VITON-HD [choi2021viton] and DressCode [morelli2022dress] at a resolution of 1024. The best and second best results are shown in bold and underline. “+R” denotes the use of reference images, and “MF” indicates mask-free inputs. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively. Unless otherwise specified, the same notations carry the same meanings throughout the figures and tables in this paper.
Method VITON-HD DressCode
[…]
Mask-based setting
RefTON (Ours) 0.079 0.870 5.96 1.05 8.91 1.15 0.056 0.899 3.28 0.76 4.84 0.83
RefTON+R (Ours) 0.072 0.873 5.25 0.97 9.10 1.41 0.052 0.902 2.84 0.65 4.73 0.76
Mask-Free setting
RefTON/MF (Ours) 0.068 0.880 5.02 0.85 8.87 1.05 0.028 0.956 1.03 0.19 4.24 0.59
RefTON+R/MF (Ours) 0.067 0.875 4.73 0.71 8.98 1.22 0.030 0.953 1.15 0.25 4.41 0.69
To further evaluate our method under high-resolution settings, Table S6 reports quantitative results on both VITON-HD and DressCode at a resolution of 1024. Across the paired and unpaired protocols, RefTON and RefTON+R consistently achieve high performance on nearly all metrics. Notably, the mask-free variants (RefTON/MF and RefTON+R/MF) deliver particularly strong results. These results demonstrate that our framework scales effectively to high-resolution synthesis and remains robust across diverse virtual try-on settings.
Appendix C Additional Qualitative Results
In this section, we provide extensive qualitative visualizations to further demonstrate the robustness, generalization ability, and high-fidelity performance of our method across different datasets, clothing categories, and evaluation settings.
We present qualitative comparisons with methods that require additional text inputs, and explicitly annotate each method’s required inputs in Fig S11(a). Our method achieves the best visual quality with the fewest required inputs. Moreover, adding a reference image further improves intricate details (e.g., lace, transparency, and texture), highlighting the advantage of visual reference. Figure S11(b) shows qualitative results on the StreetTryOn dataset under various settings. Compared with the baselines, our model produces more faithful try-on details while better preserving the person’s pose and the background. We also provide qualitative comparisons under the mask-free setting, as shown in Fig. S11(c). Our model demonstrates the strongest mask-free virtual try-on capability among all compared methods.
Figure S11: (a)We gain better results with fewer inputs required on VITON. (b)StreetTryOn qualitative results (in-the-wild & P2P): our method yields better results.(c) Our model shows stronger mask-free capability.
As shown in Fig. S12, S13, S14, S15, and S16, our approach consistently preserves garment details, structural correctness, and texture realism under both paired and unpaired scenarios, with or without mask-free (MF) inputs. Moreover, as illustrated in Fig.1, even in challenging in-the-wild conditions, our model exhibits strong robustness—maintaining accurate body pose, preserving background integrity, and producing stable try-on results without introducing artifacts or unintended changes.
Complex Patterns (VITON-HD).
As shown in Fig. S12, our model faithfully reproduces complex and fine-grained clothing patterns. Even for garments with dense textures, irregular motifs, or high-frequency visual elements, the generated results retain clear, sharp, and recognizable patterns with minimal distortion. The strong pattern-preservation ability highlights the effectiveness of our approach in capturing both global appearance and subtle local details.
Complex Structures (VITON-HD).
Fig. S13 further illustrates that our method handles garments with challenging structural designs, such as multi-layered regions, unique silhouettes, or uncommon shapes. The generated try-on results maintain correct garment geometry, coherent contours, and physically plausible spatial arrangements. This demonstrates that our framework models structural priors robustly, enabling accurate synthesis even under significant variations in shape.
DressCode Upper-body, Lower-body, and Dress Sub-sets.
As shown in Figs. S14, S15, and S16, our approach performs consistently well across the three DressCode subsets. In the unpaired and mask-free settings, our model successfully preserves fabric materials, shading, and texture characteristics while achieving realistic alignment between the garment and human body. Across diverse clothing types—including tops, pants, skirts, and full-body dresses—the synthesized results maintain stable structure, smooth boundaries, and visually coherent integration, demonstrating strong generalization and robustness.
In-the-Wild results.
In addition to controlled benchmark evaluations, RefTON demonstrates strong robustness and generalization in challenging in-the-wild scenarios. As shown in Fig. 1, it produces high-quality try-on results under diverse poses, lighting conditions, and cluttered backgrounds. Our mask-free pipeline directly transfers garments without human parsing masks or pose estimators, while preserving body pose, global structure, and identity cues. Moreover, additional-reference try-on further improves garment geometry, texture details, and overall realism. RefTON also maintains strong background consistency and avoids unnecessary changes to non-garment regions, making it reliable for real-world applications.
Figure S12: Qualitative paired results in VITON-HD dataset with complex patterns on clothes. “reference” denotes that a reference image is provided.
Figure S13: Qualitative paired results in VITON-HD dataset with complex structure on clothes. “reference” denotes that a reference image is provided.
Figure S14: Qualitative results of upper-body sub-set in Dresscode dataset unpaired setting. “reference” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.
Figure S15: Qualitative results of lower-body sub-set in Dresscode dataset unpaired setting. “reference” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.
Figure S16: Qualitative results of dresses sub-set in Dresscode dataset unpaired setting. “reference” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.